OpenStudy (anonymous):

Find the vertex of the parabola f(x)=-4x^(2)+40x-93

OpenStudy (anonymous):

I tried using the quadratic formula but im just getting (-40 plusminus sqrt112)/-8

OpenStudy (anonymous):

\[\left( -\frac{ b }{ 2a },f \left( -\frac{ b }{ 2a } \right) \right)\]

OpenStudy (anonymous):

Would I make it -40/8 or -40/-8?

OpenStudy (anonymous):

\[-\frac{ b }{ 2a }=-\left( \frac{ 40 }{ 2\cdot \left( -4 \right) } \right)=\frac{ 40 }{ 8 }=5\]

OpenStudy (anonymous):

That makes more sense. So then I'd have for my y it'd be 27?

OpenStudy (anonymous):

no

OpenStudy (marigirl):

(5,7) isn't it

OpenStudy (anonymous):

-4(5)^2+40(5)-93 = -4(25)+40(5)-93 = -100 + 200 - 93 = 7 so yes marigirl

OpenStudy (anonymous):

Oops sorry. Thanks.

OpenStudy (anonymous):

now worries rrey! you never need to apologize for effort... only lack of it!

OpenStudy (anonymous):

no, not now

OpenStudy (anonymous):

as in no worries

OpenStudy (anonymous):

as in, it's okay